# MultiVENT 2.0: A Massive Multilingual Benchmark for Event-Centric Video Retrieval

### v1

@@ Abstract
Efficiently retrieving and synthesizing information from large-scale multimodal collections has become a critical challenge. However, existing video retrieval datasets suffer from scope limitations, primarily focusing on matching descriptive but vague queries with small collections of professionally edited, English-centric videos. To address this gap, we introduce MultiVENT 2.0, a large-scale, multilingual event-centric video retrieval benchmark featuring a collection of more than 218,000 news videos and 3,906 queries targeting specific world events. These queries specifically target information found in the visual content, audio, embedded text, and text metadata of the videos, requiring systems leverage all these sources to succeed at the task. Preliminary results show that state-of-the-art vision-language models struggle significantly with this task, and while alternative approaches show promise, they are still insufficient to adequately address this problem. These findings underscore the need for more robust multimodal retrieval systems, as effective video retrieval is a crucial step towards multimodal content understanding and generation tasks.
@@ 1 Introduction
Figure 1: Example query/video pairs from MSR-VTT [38] and MultiVENT 2.0. MSR-VTT consists primarily of broad descriptive queries mapped to general video clips aimed at English-speaking audiences, while MultiVENT 2.0 targets specific current events covering 4 media formats, 6 languages, and subjects like natural disasters, politics, sports, social gatherings, and science.
@@
YouCook2 Caption Manual Cooking YouTube 15.4K 2K/15.4K 19.6s no
TVR Query Manual TV Shows TV Shows 109K 6/21.8K 76.2s no
MTVR Query Manual Movies TV Shows 218K 6/21.8K 76.2s yes
@@
MultiVENT 1.0 Caption Manual Open YouTube/Twitter 2,400 2.4K 82.6s yes
MultiVENT 2.0 (ours) Query Manual Open YouTube/Twitter 3,900 218K ? yes
Table 1: Comparison of video retrieval datasets. Many of these datasets were collected as resources for general vision-langauge model training and did not explicitly target video retrieval. Subsequent use of these collections as a retrieval benchmark leveraged descriptive video captions or subtitles as proxy queries for English only data, which differ greatly in scope and difficulty from cross-lingual text retrieval datasets [17]. To our knowledge, MultiVENT 2.0 is the first large-scale video dataset with queries explicitly developed for an event-focused multilingual retrieval task.
@@
    We introduce MultiVENT 2.0, a large-scale multilingual video retrieval task containing more than 218,000 videos. Over 3,900 queries were crafted to target aspects of current events using visual content, audio, embedded text, and text metadata from these videos.
@@ 2 Related Work
There are an increasing number of video datasets, many of which support the downstream text-video retrieval task. However, there are various limitations to these existing datasets. Many of these dataset have too few videos or have videos that are too short for comprehensive evaluation of complex event understanding [28]. For example, the popular baseline dataset, MSVD [3] only has 1,970 videos whose length average between 4 and 10 seconds. The larger and more recent VaTeX [32] and Valor-32k datasets [5] have video clips that are on average only 10 seconds long, and the clips of the extremely large dataset of over 1 million videos, HowTo100M [23], average to only 4 seconds. Many video dataset are limited to videos of a single topic or domain, such as the LSMDC dataset [26] and TVR dataset [18], which feature clips from movies and TV shows respectively, and the YouCook2 dataset which features cooking videos [42]. Thus, while these datasets are quite large in terms of number of videos, they do not represent diverse events or “real-world” topics. Moreover, while some datasets do have non-English videos, they are limited to only one non-English language; Multi-HowTo100M [14] and MTVR [19] both have only English and Chinese videos. Furthermore, despite the increase in Video datasets, very few video datasets can support event-retrieval or true multi-modal retrieval. The DiDeMo dataset [12] is a large dataset featuring event-centric videos. However, the creators of the DiDeMo dataset filtered out any edited videos, thus limiting the DiDeMo dataset to non-professional videos which thus limits the scope of the dataset. Additionally, captions or video descriptions are often extracted for pretraining as the “query” intended for the text-video retrieval task such as ActivityNet Captions [16] and MSR-VTT [38]. The V3C dataset [27] does not use captions or subtitles but video metadata. While video descriptions or summaries are high-level representations of a video, they do no reflect human style search queries befitting of a realistic retrieval task. To more easily compare and contrast the existing datasets, we have compiled the details of said video datasets in Table 1.
@@
Text-video retrieval is a core research area in video-language understanding [1, 4, 7, 31, 33, 34, 40, 41]. There have been many proposed solutions focusing on: combining pre-extracted features from frozen text and vision encoders [41]; transferring contrastive pre-training architectures, such as CLIP [24], to the video domain [10, 21, 22, 39]; leveraging video specific features such as sparsity [1, 20]; and aligning and fusing modalities [6, 5, 13, 15, 35, 37, 43]. Much emphasis has been placed on aligning visual content with text, yet videos contain more than just visual elements—they also include valuable information from text metadata, audio, and embedded text. Recently, fusion models [5, 6] and binding models [11, 43] have been developed to unify representations across modalities. For example, VAST [6] integrates vision, audio, and subtitle information for text-video retrieval, enabling higher semantic understanding of videos in a unified representation. Despite these advances, many of these models are trained on datasets not designed for single-modality retrieval, let alone retrieval across multiple modalities. As a result, these systems face challenges when handling more complex retrieval tasks requiring inference across modalities and languages. This highlights the need for more comprehensive collections to facilitate the development of more robust and adaptable video retrieval systems capable of addressing the diversity of events found in real-world multimedia content.
@@ 3 Video Collection
For MultiVENT 2.0, we extracted a large subset of videos from InternVid, filtering out those longer than five minutes. This process yielded approximately 40,000 videos for each of the five original MultiVENT 1.0 target languages: Arabic, Chinese, English, Korean, and Russian. To introduce greater variability at test time, we also included a smaller set of videos from the Spanish and ”Unknown” language categories. This ensures that systems must handle new primary languages as well as a long tail of low-resource languages.
@@ 4 Query Creation
As mentioned in Section 3, prior collections have focused more on matching descriptive aspects within videos. Accordingly, search queries have either been relatively short and vague (e.g., ”a black and white horse runs around”) or derived directly from a video’s metadata, such as a YouTube description. MultiVENT 1.0 largely followed this precedent, relying on the available metadata for each video. However, these approaches create a disconnect with modern text retrieval practices, where concise, search-engine-style English queries are increasingly used to match multilingual documents [17]. To bridge this gap, we develop a novel two-pronged approach for creating event-centric video queries: one that leverages additional fine-grained event annotations available with MultiVENT 1.0 videos [29], and another that relies solely on the videos themselves.
For this task, we recruit a team of professional linguists with expertise across the six primary languages targeted in MultiVENT 2.0 to develop the queries. Each annotator underwent training, starting with a tutorial task accompanied by detailed instructions. Afterward, we provide one-on-one feedback on their performance and closely monitor their initial annotations, offering additional guidance as needed. Finally, an annotation lead conducts a quality control check to ensure the conciseness and specificity of each query/video pair.
(a) MultiVENT Train query/video pairs by language.
(b) MultiVENT Test query/video pairs by language.
(c) MultiVENT Test query/video pairs by event type.
(d) MultiVENT Test query/video pairs by video type.
Figure 3: Breakdowns of of the number of queries mapped to relevant videos. Figure 3(a) shows that MultiVENT Train contains queries targeting the five primary languages from MultiVENT 1.0. On the other hand, in Figure 3(b) we see that MultiVENT Test adds queries targeting Spanish events to challenge systems’ multilingual robustness. As seen in Figure 3(c), Events in MultiVENT 2.0 generally map to the same categories as MultiVENT 1.0, with a long tail of infrequent event types. Figure 3(d) shows that MultiVENT 2.0 targets videos ranging from professional news broadcasts to raw first-person footage of events.
@@
While retrieving videos about overarching events is a valuable step forward, research in text information extraction (IE) has shown that individual events often encompass multiple distinct aspects, which are unlikely to be fully captured in a single piece of media. Thus, to further challenge retrieval models, annotators develop queries focused on specific, unique aspects of each event. This is facilitated by leveraging annotations from MultiVENT-Grounded [30], a collection of 1,200 MultiVENT videos containing fine-grained event annotations. These annotations, guided by adapted FrameNet event templates, include text description spans, video time intervals, and spatial bounding boxes at the frame level, and address IE-based questions such as Where did the disaster occur?, Who was affected by the disaster?, and Who responded to the disaster? Based on these annotations, annotators create 884 MultiVENT Specific queries, each targeting videos that highlight distinct aspects of the corresponding event.
@@ 4.3 Query and Video Breakdowns
To gain insight into the challenges posed by different types of video/query pairs, annotators are asked several additional questions about the videos during the query creation process. First, we ask them to confirm the video’s primary language. While InternVid ostensibly provides this information, the labels are sometimes incorrect; this is particularly common in English-labeled videos, where approximately 20% of the videos were in other languages. Figure 3(a) shows the distribution of queries targeting videos from each of the five original MultiVENT languages within MultiVENT Train, while Figure 3(b) illustrates the same for MultiVENT Test. Notably, we add Spanish as an additional primary language in the test collection to evaluate the multilingual robustness of video retrieval systems.
Additionally, annotators group events into one of the seven event types defined in MultiVENT 1.0: natural disasters (Disasters), political elections (Elections), Protests, other political developments (Political), sporting events (Sports), social events (Social Events), and scientific or technological discoveries (Science) [29]. An Other category is also provided for events that did not fit into any of these types. Figure 3(c) depicts the event type breakdown. Two notable aspects include the relative abundance of MultiVENT-specific queries for disaster events, largely due to MultiVENT-Grounded annotations prioritizing this category [30], and the predominance of Political events, which reflects the news and politics videos comprising the majority of event-based content within the InternVid collection. The Other category includes diseases, police incidents, and man-made disasters, among other infrequent event types.
@@ 5 Relevance Judgment Annotation
After the initial query creation process described in Section 4, we have a total of 6,068 annotated video/query pairs for MultiVENT-Test. In many cases, queries are only mapped to a single relevant video, and all other non-judged videos are assumed to be not relevant. This presents a challenge because, despite the targeted nature of our queries, it is likely that the distractor set contains additional relevant videos, given the size and diversity of the test collection. Furthermore, some videos may be partially relevant to a query, or relevant to some parts but not all, while others’ relevance may be unclear due to a lack of context or inherent video ambiguity.
@@
Modality Model Fine-tuned MSR-VTT MultiVENT 2.0
@@
Vision mCLIP No 0.827 0.333 0.603 0.429 0.261 0.303
OCR ICDAR $\rightarrow$ mCLIP No - 0.227 0.374 0.363 0.166 0.217
Speech Whisper $\rightarrow$ mCLIP No - 0.290 0.450 0.417 0.212 0.267
Text Description $\rightarrow$ mCLIP No - 0.293 0.491 0.445 0.228 0.284
Vision InternVideo2.0 No 0.851 0.004 0.018 0.018 0.003 0.005
All VAST No 0.739 0.118 0.118 0.198 0.080 0.116
All VAST Yes 0.896 - - - - -
All LanguageBind No 0.787 0.355 0.620 0.443 0.283 0.324
@@
To address this issue, we re-train the same pool of professional linguists from the query creation task in Section 4 to now judge the relevance of previously unseen query/video pairs. For each candidate video, annotators are asked to classify it as not relevant, possibly relevant, partially relevant, or very relevant to a query. Given the potentially vast scope of this annotation task, we limit the videos to those ranked in the top 10 by multilingual CLIP (mCLIP), a pre-trained vision-language model [2] and the strongest-performing model on MultiVENT 1.0 [29]. After removing video/query pairs already judged, we prioritize judging the highest-ranked videos, and streamline the process by grouping query/video pairs where the same video was ranked in the top 10 for multiple queries. Through this process, we collect an additional 4,396 gold relevance judgments. For evaluation purposes, we condense the middle two categories into a single category labeled somewhat relevant.
@@ 6 Baselines and Results
Jointly pre-trained vision-language models (VLMs) have recently achieved state-of-the-art results on various downstream video understanding tasks. These models typically consist of separate single-modality encoders, with a mechanism for fusing multimodal embeddings during training. In this section, we consider several prominent VLMs as baselines for video retrieval. Each of the following VLMs utilze a version of CLIP for the vision component, except for InternVid which uses a Vision Transformer (ViT) [8]:
@@
(d) Results breakdown by event type
Figure 4: Breakdown of single-modality pipeline results.
  • VALOR combines three encoders for single-modality representations with a decoder designed for multimodal text generation. [5].
  • VAST utilizes omni-modality pretraining to simultaneously fine-tune text, audio, and image encoders to enhance cross-modality learning.[6].
@@
  • Language-Bind leverages a CLIP-based encoder for all non-text modalities and has achieved state-of-the-art results on several video-understanding tasks [43]
@@
  • Whisper OCR: We utilize Whisper, a robust automatic speech recognition system, to transcribe each video’s audio track, and the resulting text is passed through mCLIP’s text encoder.
@@
Table 3 presents the performance of these baselines on MultiVENT 2.0, as well as results MSR-VTT and Valor-32k, two standard benchmark video retrieval datasets. A key observation is that while VLMs excel on prior collections, most perform poorly on our task. This is likely due to several factors: the length of the videos, as VLMs are typically trained on short video segments; and the significant domain mismatch between simple visual concepts and complex event-based natural language, the latter being much harder to map directly to visual features. This is compounded by the fact that some MultiVENT 2.0 queries specifically target non-visual aspects of the video.
Another notable takeaway is that single-modality pipeline systems show promise, though there is considerable room for improvement. Figure 4 highlights how different aspects of the problem can be better addressed by specific modalities. The distinction is particularly evident in the breakdown of video types in Figure 3(d): speech is highly effective for retrieving professional news broadcasts, yet it, along with embedded text, prove to be almost useless with raw content. In contrast, human-written descriptions and visual content are more robust across video types, with descriptions being the most useful for retrieving raw footage.

### v2

@@ Abstract
Efficiently retrieving and synthesizing information from large-scale multimodal collections has become a critical challenge. However, existing video retrieval datasets suffer from scope limitations, primarily focusing on matching descriptive but vague queries with small collections of professionally edited, English-centric videos. To address this gap, we introduce MultiVENT 2.0, a large-scale, multilingual event-centric video retrieval benchmark featuring a collection of more than 218,000 news videos and over 3,900 queries targeting specific world events. These queries specifically target information found in the visual content, audio, embedded text, and text metadata of the videos, requiring systems leverage all these sources to succeed at the task. Preliminary results show that state-of-the-art vision-language models struggle significantly with this task, and while alternative approaches show promise, they are still insufficient to adequately address this problem. These findings underscore the need for more robust multimodal retrieval systems, as effective video retrieval is a crucial step towards multimodal content understanding and generation.
@@ 1 Introduction
Figure 1: Example query/video pairs from MSR-VTT [38] and MultiVENT 2.0. MSR-VTT primarily contains broad descriptive queries mapped to general English-centric video clips, while MultiVENT 2.0 targets specific current events covering 4 media formats, 6 languages, and subjects like natural disasters, politics, sports, social gatherings, and science.
@@
YouCook2 Caption Manual Cooking YouTube 15.4K 2K/15.4K 20s no
TVR Query Manual TV Shows TV Shows 109K 6/21.8K 76s no
MTVR Query Manual Movies TV Shows 218K 6/21.8K 76s yes
@@
MultiVENT 1.0 Caption Manual Open YouTube/Twitter 2,400 2.4K 83s yes
MultiVENT 2.0 (ours) Query Manual Open YouTube/Twitter 3,900 218K 145s yes
Table 1: Comparison of video retrieval datasets. Many of these datasets were collected as resources for general vision-language model training and did not explicitly target video retrieval. Subsequent use of these collections as a retrieval benchmark leveraged descriptive video captions or subtitles as proxy queries for English only data, which differ greatly in scope and difficulty from cross-lingual text retrieval datasets [17]. To our knowledge, MultiVENT 2.0 is the first large-scale video dataset with queries explicitly developed for an event-focused multilingual retrieval task.
@@
    We introduce MultiVENT 2.0, [^2] Data available at this link. a large-scale multilingual video retrieval task containing more than 218,000 videos. Over 3,900 queries were crafted to target aspects of current events using visual content, audio, embedded text, and text metadata from these videos.
@@ 2 Related Work
There are an increasing number of video datasets, many of which support the downstream text-video retrieval task. However, there are various limitations to these existing datasets. Many of these dataset have too few videos or have videos that are too short for comprehensive evaluation of complex event understanding [28]. For example, the popular baseline dataset, MSVD [3] only has 1,970 videos whose length average between 4 and 10 seconds. The larger and more recent VaTeX [32] and Valor-32k datasets [5] have video clips that are on average only 10 seconds long, and the clips of the extremely large dataset of over 1 million videos, HowTo100M [23], average to only 4 seconds. Many video dataset are limited to videos of a single topic or domain, such as the LSMDC dataset [26] and TVR dataset [18], which feature clips from movies and TV shows respectively, and the YouCook2 dataset which features cooking videos [42]. Thus, while these datasets are quite large in terms of number of videos, they do not represent diverse events or “real-world” topics. Moreover, while some datasets do have non-English videos, they are limited to only one non-English language; Multi-HowTo100M [14] and MTVR [19] both have only English and Chinese videos. Furthermore, despite the increase in Video datasets, very few video datasets can support event-retrieval or true multi-modal retrieval. The DiDeMo dataset [12] is a large dataset featuring event-centric videos. However, the creators of the DiDeMo dataset filtered out any non-professional edited videos, thus limiting the scope of the dataset. Additionally, captions or video descriptions are often extracted for pretraining as the “query” intended for the text-video retrieval task such as ActivityNet captions [16] and MSR-VTT [38]. The V3C dataset [27] does not use captions or subtitles but video metadata. While video descriptions or summaries are high-level representations of a video, they do no reflect human style search queries befitting of a realistic retrieval task. To more easily compare and contrast the existing datasets, we have compiled the details of said video datasets in Table 1.
@@
Text-video retrieval is a core research area in video-language understanding [1, 4, 7, 31, 33, 34, 40, 41]. There have been many proposed solutions focusing on: combining pre-extracted features from frozen text and vision encoders [41]; transferring contrastive pre-training architectures, such as CLIP [24], to the video domain [10, 21, 22, 39]; leveraging video specific features such as sparsity [1, 20]; and aligning/fusing modalities [6, 5, 13, 15, 35, 37, 43]. Emphasis has been placed on aligning visual content with text, yet videos also include valuable information from text metadata, audio, and embedded text. Recently, fusion models [5, 6] and binding models [11, 43] have been developed to unify representations across modalities. VAST [6] notably integrates vision, audio, and subtitle information for text-video retrieval, enabling higher semantic video understanding in a unified representation. Despite these advances, many of these models are trained on datasets not designed for single-modality retrieval, let alone retrieval across multiple modalities. As a result, these systems face challenges when handling more complex retrieval tasks requiring inference across modalities and languages. This highlights the need for more comprehensive collections to facilitate the development of more robust and adaptable video retrieval systems capable of addressing the diversity of events found in real-world multimedia content.
@@ 3 Video Collection
For MultiVENT 2.0, we extract a large subset of videos from InternVid, filtering out those longer than five minutes. This process yields approximately 40,000 videos for each of the MultiVENT 1.0 target languages: Arabic, Chinese, English, Korean, and Russian. To introduce greater variability at test time, we also include a smaller set of videos from the Spanish and “Unknown” language categories, ensuring that systems handle both new primary languages and a long tail of low-resource languages.
@@ 4 Query Creation
As mentioned in Section 3, prior collections have focused more on matching descriptive aspects within videos. Accordingly, search queries have either been relatively short and vague (e.g., a black and white horse runs around) or derived directly from a video’s metadata, such as a YouTube description. MultiVENT 1.0 largely followed this precedent, relying on the available metadata for each video. However, these approaches create a disconnect with modern text retrieval practices, where concise, search-engine-style English queries are increasingly used to match multilingual documents [17]. To bridge this gap, we develop a novel two-pronged approach for creating event-centric video queries: one that leverages additional fine-grained event annotations available with MultiVENT 1.0 videos [29], and another that relies solely on the videos themselves.
For this task, we recruit professional linguists with expertise across the six primary languages targeted in MultiVENT 2.0 to develop the queries. Each annotator underwent training via a tutorial task with detailed instructions. Afterward, we provide one-on-one feedback on their performance and closely monitor their initial annotations, offering additional guidance as needed. Finally, an annotation lead conducts a quality control check to ensure the conciseness and specificity of each query/video pair.
Figure 3: Breakdowns of of the number of queries mapped to relevant videos. Figure 3(a) shows that MultiVENT Train contains queries targeting the five primary languages from MultiVENT 1.0. On the other hand, in Figure 3(b) we see that MultiVENT Test adds queries targeting Spanish events to challenge systems’ multilingual robustness. Figure 3(c) shows that MultiVENT 2.0 targets videos ranging from professional news broadcasts to raw first-person footage of events. Finally, as seen in Figure 3(d), events in MultiVENT 2.0 generally map to the same categories as MultiVENT 1.0, with a long tail of infrequent event types.
@@
While retrieving videos about overarching events is a valuable step forward, research in text information extraction (IE) has shown that individual events often encompass multiple distinct aspects, which are unlikely to be fully captured in a single piece of media. Thus, to further challenge retrieval models, annotators develop queries focused on specific, unique aspects of each event. This is facilitated by leveraging annotations from MultiVENT-Grounded [30], a collection of 1,200 MultiVENT 1.0 videos containing fine-grained event annotations. These annotations, guided by adapted FrameNet event templates, include text description spans, video time intervals, and spatial bounding boxes at the frame level, and address IE-based questions such as Where did the disaster occur?, Who was affected by the disaster?, and Who responded to the disaster? Based on these annotations, annotators create 884 MultiVENT Specific queries, each targeting videos that highlight distinct aspects of the corresponding event.
@@ 4.3 Query and Video Breakdowns
To gain insight into the challenges posed by different types of video/query pairs, annotators are asked several additional questions about the videos during the query creation process. First, we ask them to confirm the video’s primary language. While InternVid ostensibly provides this information, the labels are sometimes incorrect; this is particularly common in English-labeled videos, where approximately 20% of the videos were in other languages. Figure 3(a) shows the distribution of queries targeting videos from each of the five original MultiVENT 1.0 languages within MultiVENT Train, while Figure 3(b) illustrates the same for MultiVENT Test. Notably, we add Spanish as an additional primary language in the test collection to evaluate the multilingual robustness of video retrieval systems.
Additionally, linguists group events into one of the seven event types defined in MultiVENT 1.0: natural disasters (Disasters), political elections (Elections), Protests, other political developments (Political), sporting events (Sports), social events (Social Events), and scientific or technological discoveries (Science[29]. An Other category is also provided for events that did not fit into these types. Figure 3(d) depicts the event type breakdown. Two notable aspects include the relative abundance of MultiVENT-specific queries for disaster events, largely due to MultiVENT-Grounded annotations prioritizing this category [30], and the predominance of Political events, which reflects the news and politics videos comprising the majority of event-based content within the InternVid collection. The Other category includes diseases, police incidents, and man-made disasters, among other infrequent event types.
@@ 5 Relevance Judgment Annotation
After the initial query creation process, we have a total of 6,068 annotated video/query pairs for MultiVENT-Test. In many cases, queries are only mapped to a single relevant video, and all other non-judged videos are assumed to be not relevant. This presents a challenge because, despite the targeted nature of our queries, it is likely that the distractor set contains additional relevant videos, given the size and diversity of the test collection. Furthermore, some videos may be partially relevant to a query, or relevant to some parts but not all, while others’ relevance may be unclear due to a lack of context or inherent video ambiguity.
@@
To address this, we re-train the same pool of professional linguists from the query creation task in Section 4 to now judge the relevance of previously unseen query/video pairs. For each candidate video, annotators are asked to classify it as not relevant, possibly relevant, partially relevant, or very relevant to a query. Given the potentially vast scope of this annotation task, we limit the videos to those ranked in the top 10 by multilingual CLIP (mCLIP), a pre-trained vision-language model [2] and the strongest-performing model on MultiVENT 1.0 [29]. After removing video/query pairs already judged, we prioritize judging the highest-ranked videos, and streamline the process by grouping query/video pairs where the same video was ranked in the top 10 for multiple queries. Through this process, we collect an additional 4,396 gold relevance judgments. For evaluation purposes, we condense the middle two categories into a single category labeled somewhat relevant.
Modality Model MSR-VTT MultiVENT 2.0
@@
Vision mCLIP 0.827 0.333 0.603 0.429 0.261 0.303
OCR ICDAR $\rightarrow$ mCLIP - 0.227 0.374 0.363 0.166 0.217
Speech Whisper $\rightarrow$ mCLIP - 0.290 0.450 0.417 0.212 0.267
Text Description $\rightarrow$ mCLIP - 0.293 0.491 0.445 0.228 0.284
Vision InternVideo2.0 0.851 0.004 0.018 0.018 0.003 0.005
All VAST 0.739 0.118 0.118 0.198 0.080 0.116
All LanguageBind 0.787 0.355 0.620 0.443 0.283 0.324
@@ 6 Baselines and Results
Jointly pre-trained vision-language models (VLMs) have recently achieved state-of-the-art results on various downstream video understanding tasks. These models typically consist of separate single-modality encoders, with a mechanism for fusing multimodal embeddings during training. In this section, we consider several prominent VLMs as baselines for video retrieval. Each of the following VLMs utilize a version of CLIP for the vision component, with the notable exception of InternVid using a Vision Transformer (ViT) [8].
@@
Table 4: Breakdown of single-modality pipeline results. Table 4(a) shows that Chinese and Korean videos are particularly challenging, especially for vision-only models. From Table 4(b), extracted text is most beneficial for professional news broadcasts, while real-time raw footage remains challenging for all direct video content. Table 4(c) demonstrates that queries based on specific modalities are generally best handled by the corresponding modality. Finally, from Table 4(d), visual content is most effective for most MultiVENT 1.0 event types.
  • VALOR combines three encoders for single-modality representations with a decoder designed for multimodal text generation [5].
  • VAST utilizes omni-modality pretraining to simultaneously fine-tune text, audio, and image encoders to enhance cross-modality learning [6].
@@
  • Language-Bind leverages a CLIP-based encoder for all non-text modalities and has achieved state-of-the-art results on several video-understanding tasks [43].
@@
  • Whisper OCR: We utilize Whisper, a robust automatic speech recognition system [25], to transcribe each video’s audio track, and the resulting text is passed through mCLIP’s text encoder.
@@
Table 3 presents the performance of these baselines on MultiVENT 2.0, as well as results on MSR-VTT, a standard benchmark video retrieval datasets. A key observation is that while VLMs excel on prior collections, most perform poorly on our task. This is likely due to two main factors: the length of our videos, as VLMs are typically trained on short video segments; and the significant domain mismatch between simple visual concepts and complex event-based natural language, the latter being much harder to map directly to visual features. This second issue is compounded by some MultiVENT 2.0 queries targeting non-visual aspects of the video.
Another notable takeaway is that single-modality pipeline systems show promise, though there is considerable room for improvement. Table 4 highlights how different aspects of the problem can be better addressed by specific modalities. The distinction is particularly evident in the breakdown of video types in Table 4(b): speech is highly effective for retrieving professional news broadcasts, yet it, along with embedded text, prove to be almost useless with raw content. In contrast, human-written descriptions and visual content are more robust across video types, with descriptions being the most useful for retrieving raw footage.